A boat can travel on a river 10 miles upstream in the same time it takes to travel 14 miles downstream. Find the speed of the current if the speed of the boat in still water is 18 miles per hour
Answer by mananth(16949)

Boat speed = 18 mph
current speed = x mph
wind speed
speed against current 18 - x 0.07
speed with current 18 + x

Distance against wind 10 km
Distance with wind 14 km

t=d/r time against = time with
10/(18-x)= 14/(18+x)
10(18+x )=14(18 -x)
180+10x =252-14 x
14x+10 x = 72
24 x = 72
/ 24
x = 3 mph current speed